this one is a pain and I need help!
x^2+8=57
its one of these
a +-square root 65
b +-square root 57-8
c +-square root 7
d +-square root 57+8

Answer by edjones(7311)

x^2+8=57
x^2-49=0 subtract 57 from each side.
(x-7)(x+7)=0 Factor
x=7, x=-7
Ed